Solve by substitiution
4x+6y=8
y=4x-22

Answer by checkley71(8403)

Y=4X-22
4X+6(4X-22)=8
4X+24X-132=8
28X=8+132
28X=140
X=140/28
X=5 ANSWER
Y=4*5-22
Y=20-22
Y=-2 ANSWER.
PROOF
4*5+6*-2=8
20-12=8
8=8